OpenStudy (anonymous):

it a and b are roots of equation 3x^2 - x + 8=0 the equation whose roots are 1/(a+2) and 1/(b+2) is?

OpenStudy (shubhamsrg):

let second eqn be x^2 + px + q =0 for the second eqn,, sum of roots = 1/(a+2) + 1/(b+2) = (a+b)+4/ (ab + 2(a+b) +4) = -p product of roots = 1/a+2 * 1/b+2 = 1/ (ab + 2(a+b) +4) =q from 1st given eqn,,we know a+b = 1/3 and ab =8/3 just substitute these values to find p and q and thus the req eqn note:alternative method would be to find a and b and then using it find 2nd eqn.

OpenStudy (anonymous):

@satellite73

OpenStudy (shubhamsrg):

ohh lol.. :|

OpenStudy (anonymous):

di nt understand

OpenStudy (anonymous):

a +b = -b/a that method lol

OpenStudy (shubhamsrg):

sum of roots things..hmm..

OpenStudy (anonymous):

yuo

OpenStudy (anonymous):

can u solve by that

OpenStudy (shubhamsrg):

i used that method only buddy!! hehehe alternative which i suggested was some other method :D

OpenStudy (anonymous):

thxx